A manufacturer at present has 2500 units of product in stock. The product is now selling at $4 per unit. Next month the unit price will increase by $0.50. The manufacturer wants the total revenue received from the sale of the 2500 units to be less than $10750. What is the maximum number of the units that can be sold this month?
Answer by FrankM(1040)

EQ1 X+Y=2500
EQ2 4X+4.5Y=10750
Rearrange EQ1 to solve for Y in terms of X Y=2500-X
4X+4.5(2500-X) = 10750
4X -4.5X + 11250 = 10750
-.5X= -500
X= 1000
1000 units at $4 plus 1500 units at $4.5 add to exactly $10,750